# L-WISE: Boosting human image category learning through model-based image selection and enhancement

### v1

@@ 3.1 Training task-specific robustified models
To obtain robustified models for a task specific category space, we adversarially-trained ResNet-50 ANNs (He et al., 2016) on the ImageNet-1K (Deng et al., 2009) and iNaturalist 2021 (Van Horn et al., 2021) datasets (separately) using the same technique from Mądry et al. (2018). To adapt the resulting model to the three categorization tasks of interest, we conducted additional adversarial fine-tuning on the smaller datasets: a small subset of moth species images from iNaturalist, the HAM10000 skin lesion dermoscopy dataset (Tschandl et al., 2018), and the MHIST colon histology dataset (Wei et al., 2021).
@@ 3.2 Predicting category recognition difficulty
We propose an elegant way to predict the human categorization error rate on a given image, which suggests a new image-recognition difficulty score: the logit activation (pre-softmax) at the ground truth category output. The higher this logit value is, the lower the human categorization error rate. We established this relationship through user study reports in a basic natural image categorization task with 16 animal categories (Fig. 1A). We found this robustified model-based metric to be the current state-of-the-art in predicting human error rates (see Appendix Fig. S4).
3.3 Generating image perturbations to enhance category percept
Given a bona fide metric of image difficulty, backpropagation on the model from output back to pixel space allows us to perturb an image to be easier to recognize with respect to its ground truth label. To this end, we impose a criterion to maximize the ground-truth logit activation, while also constraining the perturbations to be within a predefined $\ell_{2}$-norm pixel-budget. This approach is analogous to  Gaziv et al. (2024) but is designed to enhance the ground truth percept, rather than guiding away from it.
@@ 4 Results
Figure 3: Novice learners who had their curriculum augmented by our method showed improved test-time categorization accuracy for previously unfamiliar categories. Empirical results for a 4-way fine-grained moth species categorization task. Panel A shows examples of the four moth classes in the task, side-by-side with their model-based enhanced versions at the highest pixel budget used in our experiments ($\epsilon=8$). While subtle, one notable difference is the distinctive wing spots of moth class 2, which are enlarged in the enhanced version of the image. Also shown are difference images showing the (5x magnified) difference between original and enhanced versions, and heat maps with more red coloration in regions of larger changes from enhancement. B compares the average smoothed accuracy of participants in the L-WISE group and a control group. Shaded areas denote 95% CI by bootstrap. The test accuracy gain of the L-WISE group relative to the control group is statistically significant (chi-square, $p<0.001$). C,D show the trial-dependent empirical profiles of the average image difficulty percentile of selected images, which (noisily) step-wise increases, and the perturbation pixel budget for enhancement ($\epsilon$), which step-wise decreases. The profiles of the baseline controls (black dotted lines) are uniform, indicating randomly-selected images with no enhancement.
@@
Figure 4: Our approach can boost accuracy and time efficiency of image category learning for humans across varied image domains, including clinically-relevant ones. Panel A compares the mean test-time accuracy and training-phase duration of human participants who learned a moth photo, dermoscopy, or histology classification task, randomized to L-WISE and control groups. All differences between L-WISE and the control group are statistically significant (chi-squared test, $p<0.05$). Panel B shows how precision and recall in L-WISE and control groups, with each point representing a specific class in one of the three tasks. All error bars show 95% bootstrap confidence intervals. Each class from the dermoscopy and histology tasks is illustrated in panels C and D respectively, similarly to the moth classes in Fig. 3A.
@@
We used robustified ANNs to both enhance images and predict the difficulty of images across multiple domains. We applied each of these techniques to improve the final test performance (on original, unmodified images) of novices learning challenging image classification tasks.
@@
4.1 Models can both predict image recognition difficulty and reduce it
@@
We tested the effects of image enhancement via maximization of the ground truth logit from robustified ANNs on human image categorization accuracy. We successfully demonstrate that we can enhance images by maximizing the ground truth logit from a robustified ANN (ResNet-50) using gradient descent in image pixel space. As the size of the perturbations grows ($\ell_{2}$-norm pixel budget $\epsilon$), human participants become increasingly accurate on a 16-way animal photograph classification task derived from ImageNet (Fig. 1B1, chance = 1/16). While mean accuracy on the original, unmodified ($\epsilon=0$) images was 0.75, mean accuracy on enhanced images was as high as 0.84 at $\epsilon=20$. The accuracy gains from enhancement appear to reach a saturation point as the perturbations grow larger. The improvements in accuracy are also somewhat dependent on the starting ground truth logit score, as shown in Appendix Fig. S8: accuracy gains are significantly higher for “difficult” images than for “easy” images. Baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE), Multi-Scale Retinex with Color Restoration (MSRCR), and Adobe Photoshop Lightroom’s “Auto” enhancement feature (LR) had no significant effect on performance, despite inducing image perturbations of considerably larger $\ell^{2}$ norm on average than the $\ell^{2}$ pixel budget $\epsilon$ values we used for model-based enhancement.
We also demonstrate that the robustified model’s ground truth logit $L_{\text{gt}}(x)$ is strongly correlated with the rate at which humans choose the ground truth category associated with image $x$ in a 16-way basic animal classification task (Fig. 1A). We used robustified ResNet-50 to calculate $L_{\text{gt}}$ for each of the 2,400 distinct natural images used in the task, and applied logistic regression to predict binary correct vs. incorrect responses to individual image trials (pooling responses to original images and MSRCR, CLAHE, and LR control images to increase sample size). The logistic regression model (Fig. 1A) used $L_{\text{gt}}$ to predict the correctness of the trial responses with $\textbf{AUC}=0.71$ (10-fold cross-validation, $p<0.001$). Notably, we found this simple approach to be better aligned with human error rates than the c-score, a state-of-the-art metric at predicting the difficulty of images for humans (Mayo et al., 2023) (see Appendix Fig. S4). We also demonstrate difficulty prediction and image enhancement with XCiT vision transformers (see Appendix Figs. S5-S7).
@@ 4.2 L-WISE improves both learning speed and test accuracy for humans
We applied both image difficulty prediction and image enhancement in a novel framework that designs curriculum image sequences for novices learning challenging image classification tasks. Our Logit-Weighted Image Selection and Enhancement (L-WISE) algorithm operates on image trial sequences used for training human participants on image classification tasks. The performance of each participant is subsequently evaluated in a testing phase, which includes randomly-selected, unmodified images (unaffected by L-WISE). During the early portion of the training phase, L-WISE randomly selects images from below a certain difficulty percentile that linearly increases as the training phase progresses. Selected images are enhanced at each trial during this period, within an $\ell_{2}$ pixel budget $\epsilon$ that decreases in a stepwise-exponential fashion (Fig. 3C-D).
We tested L-WISE’s efficacy at improving test-time accuracy of human learners in three challenging image category learning tasks (Figs. 3-4). Participants were randomly assigned to an L-WISE group or a control group (randomly-selected, non-enhanced images). L-WISE increased the average test-time accuracy margin above chance levels by 57.6% on a 4-way moth classification task ($p<0.001$ on chi-squared test), by 72.3% on a 4-way skin lesion dermoscopy task ($p<0.001$), and by 33.1% on a binary colon polyp histology task ($p=0.023$) (Fig. 4A). In all three tasks, participant accuracy in the L-WISE group increased initially and then declined to varying degrees as more and more difficult images were selected and the degree of enhancement simultaneously reduced. In addition to improving test-time accuracy, L-WISE decreased the mean time to learn the task (with a fixed number of training trials) by 20% for the moth task, 23% for the dermoscopy task, and 22% for the histology task (Fig. 4A).
@@ 4.3 Image enhancement and selection both contribute to efficacy of L-WISE
Mean Acc. Training Time Mean Acc. Training Time
@@
Table 1: Both image enhancement tapering (ET) and image difficulty selection (DS) contribute to ability of L-WISE to assist learners. The benefits of image enhancement are dependent on easy-to-hard sequencing (“ET” outperforms “ET (shuffled)”), but the benefits of difficulty-based selection appear to stem from simply showing an easier distribution of images during training (“DS (shuffled)” is as good as or better than “DS”). Training times are in minutes. Parentheses show 95% bootstrap confidence intervals for the mean with 10,000 bootstrap replicates. * denotes a significant difference in accuracy from the control group ($p<0.01$, $\chi^{2}(1)$ test).
We tested several ablated versions of L-WISE to determine the relative contributions and potential additive benefits of its components, which include (A) image enhancement based on logit optimization, (B) image selection based on logit-estimated difficulty, and (3) easy-to-hard curriculum trends enabled by A and B. The results of our ablation study, which employed the moth classification and skin lesion dermoscopy tasks, are shown in Table 1. In “Enhancement Tapering,” (ET) only the enhancement component of L-WISE is used, with no difficulty-based image selection. Conversely, in “Difficulty Selection” (DS), images are not enhanced. In “ET (shuffled)” and “DS (shuffled),” the order of training trials is randomly permuted after either the ET or DS intervention. We hypothesized that shuffling would abolish at least some of the potential accuracy gains attributable to ET and DS due to the possible usefulness of easy-to-hard ordering.
The results show that both ET and DS have strong benefits in isolation. ET increased the test-phase accuracy margin above chance by 46.8% for the moth task and 56.5% for the dermoscopy task, while DS increased the same margin by 8.1% (not significant) and 53.2% respectively. ET (shuffled) was less effective, increasing the margin above chance by 23.0% for the moth task and 11.2% (not significant) for the dermoscopy task. Surprisingly, DS (shuffle) outperformed DS without shuffling, increasing the margin above chance by 45.2% for the moth task and 58.2% for the dermoscopy task (the increase of DS (shuffle) relative to DS is statistically significant for the moth task only). Unablated L-WISE numerically outperformed all ablated conditions, increasing the margin above chance by 57.6% for the moth task and 72.3% in the dermoscopy task. However, additional paired comparisons indicated that the differences between these increases and those from ET or DS (shuffled) are not statistically significant for either task. ET and DS did demonstrate a statistically significant additive benefit in terms of learning efficiency, however. On the moth task, training time was 6% shorter for L-WISE than the next fastest group, which was ET. Similarly, on the dermoscopy task, training time was 5% shorter for L-WISE than the next fastest group, DS (shuffled).
@@ 5 Discussion
In this study, we demonstrate that the leading models of the ventral visual stream, the robustified ANNs, can be used to both predict the empirical image recognition difficulty as reported by humans, and generate “perturbed”, enhanced, versions of it that are easier to categorize as the ground truth. We harness these properties to design a model-based curriculum design for human image category learning. We show that a combination of selecting images within a certain difficulty range and perturbing those images to enhance the perception of the ground truth category leads to substantial improvements in human test-time classification performance under realistic conditions.
The results of our ablation study show that at least a portion of these improvements can be achieved with image enhancement alone: humans can learn from distorted images and subsequently achieve superior generalization to unseen examples (Table 1). There are several possible explanations for this effect. Image enhancements might draw the learner’s attention to relevant features, such as the distinctive dot in the middle of each wing of the Idaea biselata moth in Fig. 3B (Hufnagel, 1767), or the multiple colors and irregular borders that appear to be enhanced in the melanoma image of Fig. 4 (Tsao et al., 2015). Enhancements might also attenuate features that distract from or contradict the ground truth: for example, in Appendix Fig. S8B, buffalo standing behind the ground truth “antelope” are variously blurred or nearly erased. Analogously, images with high ground truth logits (low predicted difficulty) might tend to have clear class-relevant features and few distracting or contradictory features. These parallel explanations for the effects of image enhancement and image selection could help explain why the “enhancement tapering” and “difficulty selection” strategies in isolation provide comparable accuracy gains to each other, and why combining both strategies did not lead to large additive improvements in accuracy (although full L-WISE did enable significantly shorter training time than either strategy alone).
@@ 5.1 Limitations
This work is a proof-of-concept demonstration that robustified models can be applied to augment image category learning in humans. We did not exhaustively search for optimal curriculum design strategies image enhancement hyperparameters, nor did we study the “dose-dependency” of image enhancement or selection. L-WISE applies a fixed schedule for all learners: human learning could plausibly be boosted further by adapting the degree of image enhancement or the difficulty of selected images to the learner’s progress in real time (Lu & Dosher, 2022; Mettler & Kellman, 2014).
One caveat to our approach is that logit maximization can sometimes appear to have a homogenizing effect on image distributions. For example, “benign mole” dermoscopy images enhanced with high $\epsilon$ budgets tend to all resemble smooth and uniform blobs (see appendix Fig. S9H for an example): this clearly illustrates a task-relevant difference from melanoma (which tends to be asymmetric with irregular borders (Tsao et al., 2015)), but obscures much of the real-world heterogeneity among benign moles. A similar risk might apply to image selection: images with high ground truth logits might belong to limited regions of the overall class distribution. Biased perturbations or selections reflecting biases in the underlying datasets are another concerning possibility, particularly for dermoscopy (Daneshjou et al., 2022). These caveats must be thoroughly investigated before real-world educational applications of L-WISE, particularly in the medical domain.

### v2

@@ 3.1 Training task-specific robustified models
To obtain robustified models for a task specific category space, we adversarially trained ResNet-50 ANNs (He et al., 2016) on the ImageNet-1K (Deng et al., 2009) and iNaturalist 2021 (Van Horn et al., 2021) datasets (separately) using the same techniques as Mądry et al. (2018). To adapt the resulting model to the three categorization tasks of interest, we conducted additional adversarial fine-tuning on the corresponding smaller datasets: a subset of moth species images from iNaturalist (from iNaturalist pretraining), the HAM10000 dermoscopy dataset (Tschandl et al., 2018) (ImageNet pretraining), and the MHIST histology dataset (Wei et al., 2021) (ImageNet pretraining).
@@ 3.2 Predicting category recognition difficulty
We propose a simple approach to predicting the human categorization error rate on specific images, in the form of a new image recognition difficulty score: the logit activation (pre-softmax) at the ground truth category output unit of a robustified ANN. The higher this logit value is, the lower the human categorization error rate. We establish this relationship through human participant responses during a natural image categorization task with 16 basic animal categories (Fig. 1A). We found this logit score to be the new state-of-the-art in predicting human error rates (see Appendix Fig. S6).
3.3 Generating image perturbations to enhance category percepts
While the ground truth logit score predicts image difficulty at baseline, we show that we can also generate perturbations that maximize this value by backpropagating from the logit score to pixel space and running projected gradient ascent to update the pixel values, limiting the “pixel budget” $\ell_{2}$ norm of the perturbation to a value $\epsilon$. These perturbations make images easier for humans to recognize with respect to their ground truth labels (Fig. 1B). This approach is analogous to  Gaziv et al. (2023), but is designed to enhance the ground truth percept rather than guiding away from it.
3.4 Boosting image category learning in humans
Our approach to augmenting visual learning is summarized in Fig. 2. In the naive baseline scenario (Fig. 2A-C), the novice human learner is presented with a sequence of training trials through an online platform. In each training trial, a randomly selected image from one of $N$ categories is presented, and the learner attempts to choose the correct category among $N$ possible labels. The learner receives feedback after each training trial indicating the correct category label and whether or not their response was correct. In most of our experiments, to ensure that participants begin at the chance level and to avoid possible priors induced by the category names, we assign each category to a random Greek name unrelated to the task (with random assignments for each participant). After completion of the training phase, the experiment transitions into a test phase, in which the same task continues over a held-out set of images and no feedback is provided (Fig. 2C). During the test phase we measure the main visual learning outcome of interest, the test accuracy.
Harnessing key observations of robustified ANNs, our “L-WISE” approach uses one such model to optimize the training phase in a way that improves the test accuracy via two mechanisms: (i) sampling training images based on their predicted recognition difficulty, and (ii) enhancing the training images. We adjust the “strength” of both mechanisms in a time-dependent manner: for the former, we set the maximum allowable difficulty score for an image to be shown in a given trial, and for the latter, we limit the enhancement perturbation magnitude to an $\ell_{2}$ norm pixel-budget $\epsilon$. The user of our approach can flexibly define arbitrary time-dependent profiles for image selection and enhancement (Fig. 2D-F). In this study, we focused on a linear ramp schedule for the allowable image difficulty at a given time, and on exponential tapering of the enhancement $\epsilon$. Intuitively, this should correspond to an easy-to-challenging trend during the training phase. Extensively optimizing these schedules was not a goal of our study, which rather serves as a proof-of-concept. Incorporating image selection and enhancement led to significant gains in the accuracy of human participants on unmodified, randomly-selected test images, while also reducing the time needed to complete the training phase (which consists of a constant number of trials). This result was robustly observable across the varied image domains and category spaces we tested.
@@ 4 Results
Figure 3: Novice learners who had their curriculum augmented by our method showed improved test-time categorization accuracy for previously unfamiliar categories. This figure shows empirical results for a 4-way fine-grained moth species classification task. Panel A shows examples of the four moth classes, side-by-side with their model-enhanced versions at the highest pixel budget used in our experiments ($\epsilon=8$). While subtle, one notable difference is the distinctive wing spots of moth class 2, which are enlarged in the enhanced version of the image. Also included are difference images showing the (5x magnified) difference between original and enhanced versions, and heat maps with more red coloration in regions of larger changes from enhancement. B compares the average smoothed accuracy of participants in the L-WISE group and a control group. Shaded areas denote 95% CI by bootstrap. The test accuracy gain of the L-WISE group relative to the control group is statistically significant ($\chi^{2}(1)$ test, $p<0.001$). C,D show the trial-dependent empirical profiles of the average image difficulty percentile of selected images, which (noisily) increases step-wise, and the perturbation pixel budget for enhancement ($\epsilon$), which decreases step-wise. The profiles of the baseline controls (black dotted lines) are uniform, indicating randomly-selected images with no enhancement.
@@
Figure 4: Our approach can boost time efficiency and final accuracy of image category learning for humans across varied image domains, including in clinically relevant tasks. Panel A compares the mean test-phase accuracy and training-phase duration of human participants who were randomized to L-WISE or control groups and learned a moth photo, dermoscopy, or histology classification task. All differences between L-WISE and the control group are statistically significant ($\chi^{2}(1)$ test, $p<0.05$). Panel B shows precision and recall in L-WISE and control groups, with each point representing a specific class in one of the three tasks. All error bars show 95% bootstrap confidence intervals. Each class from the dermoscopy and histology tasks is illustrated in panels C and D respectively, similarly to the moth classes in Fig. 3A.
@@
We used robustified ANNs to both enhance images and predict the difficulty of images across multiple domains. We applied both of these techniques to improve the final test performance (on original, unmodified images) of novices learning challenging image classification tasks.
@@
4.1 Robust models can both predict image recognition difficulty and reduce it
@@
We tested the effects of a novel model-based image enhancement algorithm on human image categorization accuracy. We demonstrate that we can enhance images by maximizing the ground truth logit from a robustified ANN (ResNet-50) using gradient ascent in image pixel space. As the magnitude of the enhancement perturbations grows ($\ell_{2}$ norm pixel budget $\epsilon$), human participants become increasingly accurate on a 16-way animal photograph classification task derived from ImageNet (Fig. 1B, chance = 1/16). While mean accuracy on the original, unmodified ($\epsilon=0$) images was 0.75, mean accuracy on enhanced images was as high as 0.84 at $\epsilon=20$. The accuracy gains from enhancement appear to approach a saturation point as the perturbations grow larger. The improvements in accuracy are also somewhat dependent on the starting ground truth logit score, as shown in Appendix Fig. S11: accuracy gains are larger for “difficult” images than for “easy” images. Baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE, Zuiderveld (1994)), Multi-Scale Retinex with Color Restoration (MSRCR, Jobson et al. (1997); Petro et al. (2014)), and Adobe Photoshop Lightroom’s “Auto” enhancement feature (LR, Adobe Inc. (2024)) had no significant effect on human accuracy, despite inducing image perturbations of considerably larger $\ell^{2}$ norm on average than the $\ell^{2}$ pixel budget $\epsilon$ values we used for model-based enhancement.
[…]
@@ 4.2 L-WISE improves both learning speed and test accuracy for humans
We applied both image difficulty prediction and image enhancement as part of a novel method that designs curriculum image sequences for novices learning challenging image classification tasks. Our proposed algorithm, “Logit-Weighted Image Selection and Enhancement” (L-WISE), operates on image trial sequences used to train human participants on image classification tasks through trial-by-trial feedback. The performance of each participant is evaluated in a subsequent testing phase without feedback, which includes only randomly-selected, unmodified images (unaffected by L-WISE). During the early portion of the training phase, L-WISE randomly selects images from below a certain difficulty percentile that linearly increases as the training phase progresses. Selected images are enhanced at each trial during this period, within an $\ell_{2}$ pixel budget $\epsilon$ that decreases in a stepwise-exponential fashion (Figs. 2E-F and 3C-D).
We tested L-WISE’s efficacy at improving test-time accuracy and training duration of human learners on three challenging image category learning tasks (Figs. 3-4). Participants were randomly assigned to a control group with randomly-selected, non-enhanced images throughout the task, or to an L-WISE group. L-WISE increased the average test-time accuracy margin above chance levels by 57.6% on a 4-way moth species classification task ($p<0.001$ on $\chi^{2}(1)$ test), by 72.3% on a 4-way skin lesion dermoscopy task ($p<0.001$), and by 33.1% on a binary colon histology task ($p=0.023$) (Fig. 4A). In all three tasks, participant accuracy in the L-WISE group increased initially and then declined to varying degrees as more and more difficult images were selected and the degree of enhancement was simultaneously reduced. In addition to improving test-time accuracy, L-WISE decreased the mean time to learn the task (with a fixed number of training trials) by 20% for the moth task, 23% for the dermoscopy task, and 22% for the histology task (Fig. 4A).
@@ 4.3 Image enhancement and selection both contribute to efficacy of L-WISE
Mean acc. Training duration Mean acc. Training duration
@@
Table 1: Both image enhancement tapering (ET) and image difficulty selection (DS) contribute to the ability of L-WISE to assist learners. The benefits of image enhancement are dependent on easy-to-hard sequencing (“ET” outperforms “ET (shuffled)”), but the benefits of difficulty-based selection appear to stem from simply showing an easier distribution of images during training (“DS (shuffled)” performs as well as or better than “DS”). Training durations are in minutes. Parentheses show 95% bootstrap confidence intervals for the mean with 10,000 bootstrap replicates. * denotes a significant difference in accuracy from the control group ($p<0.01$, $\chi^{2}(1)$ test).
We tested several ablated versions of L-WISE to determine the relative contributions of its components, which include (A) image enhancement based on logit maximization, (B) selection of relatively easy images according to logit-estimated difficulty, and (3) easy-to-hard curriculum trends enabled by A and B. The results of our ablation study, which employed the moth classification and skin lesion dermoscopy tasks, are shown in Table 1. In “Enhancement Tapering,” (ET) only the enhancement component of L-WISE is used, with no difficulty-based image selection. Conversely, in “Difficulty Selection” (DS), images are selected based on difficulty but not enhanced. In “ET (shuffled)” and “DS (shuffled),” the order of training trials is randomly permuted after the ET or DS intervention respectively. We hypothesized that shuffling would abolish at least some of the potential accuracy gains attributable to ET and DS due to the possible helpfulness of easy-to-hard trends.
The results show that both ET and DS have significant benefits in isolation. ET increased the test-phase accuracy margin above chance by 46.8% for the moth task and 56.5% for the dermoscopy task, while DS increased the same margin by 8.1% (not significant) and 53.2% respectively. ET (shuffled) was less effective, increasing the margin above chance by 23.0% for the moth task and 11.2% (not significant) for the dermoscopy task. Surprisingly, DS (shuffled) outperformed DS without shuffling, increasing the margin above chance by 45.2% for the moth task and 58.2% for the dermoscopy task (the increase of DS (shuffled) relative to DS is statistically significant for the moth task only). Unablated L-WISE numerically outperformed all ablated conditions, increasing the margin above chance by 57.6% for the moth task and 72.3% in the dermoscopy task. However, additional paired comparisons indicated that the differences between these increases and those from ET or DS (shuffled) are not statistically significant for either task. ET and DS did demonstrate a statistically significant additive benefit in terms of learning speed, however. On the moth task, training duration was 6% shorter for L-WISE than for the next fastest group, which was ET. Similarly, on the dermoscopy task, training duration was 5% shorter for L-WISE than for the next fastest group, DS (shuffled).
@@ 5 Discussion
In this study, we demonstrate that robustified ANNs can be used to both predict the empirical recognition difficulty of individual images as reported by humans, and also generate enhanced versions of images that are easier for humans to correctly categorize. We harness these capabilities to develop a model-based curriculum design algorithm to augment human image category learning. We show that a combination of selecting images within a certain difficulty range, and perturbing those images to enhance the perception of the ground truth category, leads to substantial improvements in human training speed and test-time classification performance on randomly-selected, unmodified images.
The results of our ablation study show that at least a portion of these improvements can be achieved with image enhancement alone: humans can learn from perturbed images and subsequently achieve superior generalization to unseen examples (Table 1). There are several possible explanations for this effect. Image enhancements might draw the learner’s attention to relevant features, such as the distinctive dot in the middle of each wing of the Idaea biselata moth in Fig. 3B (Hufnagel, 1767), or the irregular border and multiple colors that appear to be enhanced in the melanoma image in Fig. 4C (Tsao et al., 2015). Enhancements might also diminish features that distract from or contradict the ground truth: for example, in Appendix Fig. S11B (second image from the left), buffalo standing behind the ground truth “antelope” are variously blurred or nearly erased. Analogously, images with high ground truth logits (low predicted difficulty) might tend to have clearer class-relevant features and fewer distracting or contradictory features. These parallel explanations for the effects of image enhancement and image selection could help clarify why the “enhancement tapering” and “difficulty selection” strategies in isolation provide comparable accuracy gains to each other, and why combining both strategies did not lead to large additive improvements in accuracy (although additive increases in learning speed were observed).
@@ 5.1 Limitations
This work is a proof-of-concept demonstration that robustified ANNs can be applied to augment image category learning in humans. We did not exhaustively search for optimal curriculum design strategies or image enhancement hyperparameters, nor did we study the “dose-dependency” of image enhancement or selection. L-WISE applies a fixed schedule of maximal image difficulty and image enhancement magnitude for all learners: human learning could plausibly be augmented more effectively by adapting the degree of image enhancement and the difficulty of selected images to the learner’s progress in real time (Lu & Dosher, 2022; Mettler & Kellman, 2014).
One caveat to our approach is that logit maximization can sometimes appear to have a homogenizing effect on image distributions. For example, “benign mole” dermoscopy images enhanced with high $\epsilon$ budgets tend to all resemble smooth and uniform blobs (see Appendix Fig. S12H for an example). This clearly illustrates a task-relevant difference from melanoma (which tends to be asymmetric with irregular borders (Tsao et al., 2015)), but obscures much of the real-world heterogeneity among benign moles. A similar risk might apply to image selection: images with high ground truth logits might belong to limited regions of the overall class distribution. Biased perturbations or selections reflecting biases in the underlying datasets are another concerning possibility, particularly for dermoscopy (Daneshjou et al., 2022). These caveats must be thoroughly investigated before deployment of real-world educational applications of L-WISE, especially for clinical tasks with potential patient safety ramifications.